OpenStudy (anonymous):

Find the inverse function of f. f(x)=4-5x^3

OpenStudy (anonymous):

can you solve \[-131=4-5\times x^3\] for \(x\)?

OpenStudy (anonymous):

f^-1(x)=^3sqrt 4/5-1/5x

OpenStudy (anonymous):

x=4-5y^3 solve for y

OpenStudy (anonymous):

\[ x=4-5[f^{-1}(x)]^3 \]

OpenStudy (anonymous):

im very confused

OpenStudy (anonymous):

i will ask the question again if it is not clear how to do it, then we need to do this one first solve \[-132=4-5x^3\]

OpenStudy (anonymous):

yeah 5 of us are giving different approaches

OpenStudy (anonymous):

oops i meant \[-131=4-5x^3\] solve for \(x\) in 3 steps

OpenStudy (anonymous):

I didn't tell you how to do it, I just gave you the answer.

OpenStudy (anonymous):

and lets keep track of the steps when we do it

OpenStudy (anonymous):

i need to know how to do it

OpenStudy (shamil98):

Inverse functions are basically that function the opposite flow for example if F(x) = 2x +3 f'(x) = (y-3)/2

OpenStudy (anonymous):

backing out of the crowded room

OpenStudy (shamil98):

f^-1*

OpenStudy (shamil98):

I still have calculus in ma head ._.

OpenStudy (anonymous):

\[-131=4-5x^3\] subtract \(4\) from both sides, get \[-135=-5x\] divide both sides by \(-5\) \[-27=x^3\] take the cubed root of both sides \[\sqrt[3]{27}=x\] or \[x=3\]

OpenStudy (anonymous):

where did 131 come from?

OpenStudy (anonymous):

now we can do the same thing to solve \[x=4-5y^3\] but this time without numbers subtract \(4\) from both sides \[x-4=-5y^3\] divide both sides by \(-5\) \[\frac{x-4}{-5}=y^3\] take the cubed root of both sides \[\sqrt[3]\frac{x-4}{-5}=y\] or \[y=-\sqrt[3]{\frac{x-4}{5}}\]

OpenStudy (anonymous):

i made up the \(-131\) because if you can do it with \(-131\) you can to it with \(x\) but if you cannot, then you cannot

OpenStudy (anonymous):

Oh I get it

OpenStudy (anonymous):

find the inverse of x/x+9

OpenStudy (anonymous):

once again, how would you solve \[7=\frac{x}{x+9}\]?

OpenStudy (anonymous):

No idea

OpenStudy (anonymous):

i would multiply by \(x+9\) on both sides first

OpenStudy (anonymous):

\[7=\frac{x}{x+9}\\7(x+9)=x\]

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

distribute

OpenStudy (anonymous):

right \[7x+7\times 9=x\] then what?

OpenStudy (anonymous):

7x+63=x

OpenStudy (anonymous):

after that i mean

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

you want to get \(x\) on one side of the equal sign at the moment you have \(7x\) on the left, and \(x\) on the right subtract \(x\) from both sides

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

what is \(7x-x\)?

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

ok good and what is \(x-x\)?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

right, so you have \[6x+63=0\] now subtract \(63\) from both sides

OpenStudy (anonymous):

x=10.5

OpenStudy (anonymous):

ok good now lets review the steps, because we are going to have to do them again, but without numbers 1) multiply both sides by \(x-9\) 2) distribute 3) subtract \(x\) from both sides 4) put everything without \(x\) on the other side 5) divide by the coefficient of \(x\)

OpenStudy (anonymous):

now we solve \[x=\frac{y}{y+9}\] for \(y\) 1) multiply both sides by \(y+9\) and get \[x(y+9)=y\] 2) distribute \[xy+9x=y\] 3) subtract\(y\) from both sides \[xy-y+9x=0\] 4) subtract \(9x\) from both sides \[xy-y=-9x\]

OpenStudy (anonymous):

now we have a step we did not do before we have to factor out the \(y\) on the left, because we don't know what \(xy-y\) is unlike say \(7y-y\) which we know is \(6y\) here you get \[(x-1)y=-9x\]

OpenStudy (anonymous):

last step is to divide, giving \[y=\frac{-9x}{x-1}\]